# Simplex volumes in hyperplane arrangements

### v1

@@ Abstract
The Erdős’s distinct distance problem and unit distance problem are classical problems posed by Erdős [12] in 1946, which ask for the minimum number of distinct distances determined by $n$ points in $\mathbb{R}^{d}$ and the maximum number of pairs at a given distance that such a set can form, respectively. These problems have been studied for many decades, and a variety of interesting results have been obtained. Well-studied variants of these problems ask for the minimum number of distinct volumes of $d$-simplices determined by $d+1$ points among $n$ points in $\mathbb{R}^{d}$, and the maximum number of $d$-simplices with unit volume among $n$ points in $\mathbb{R}^{d}$. We propose the following dual setting of these problems and give partial results:
  • Determine the maximum number of unit-volume $d$-simplices determined by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number of minimum/maximum-volume $d$-simplices determined by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct volumes.
@@ 1 Introduction
For $d\geq 2$, a set of points $X$ in $\mathbb{R}^{d}$ with $|X|\geq d+1$ is said to be in general position if no $d+1$ points from $X$ lie on the same (d1)(d-1)-dimensional hyperplane, and we say that an arrangement of hyperplanes $\mathcal{A}$ in $\mathbb{R}^{d}$ with $|\mathcal{A}|\geq d+1$ is in general position if every $d$ hyperplanes intersect at a single point and no $d+1$ have a common point.
Throughout this paper, we use $O_{d}$ to denote the standard big-$O$ notation where the constant may depend on $d$. We adopt similar conventions for $\Omega_{d}$ and $\Theta_{d}$.
1.1 Erdős’s distinct distance problem
In 1946, Erdős [12] posed the question: what is the minimum number $g_{d}(n)$ of distinct distances determined by $n$ points in $\mathbb{R}^{d}$. This question is known as Erdős’s distinct distance problem, and a variety of interesting results have been obtained. For the planar case $d=2$, Erdős showed an upper bound $g_{2}(n)=O(n/\sqrt{\log n})$ by considering the $\sqrt{n}\times\sqrt{n}$ square lattice. A lower bound for $g_{2}(n)$ has been improved many times [31, 3, 4, 40, 37, 43, 27]. The most recent bound is due to Guth and Katz [25], who showed that $g_{2}(n)=\Omega(n/\log n)$, which almost matches Erdős’s upper bound.
In higher dimensions, considering a $d$-dimensional grid $n^{1/d}\times\cdots\times n^{1/d}$, we obtain $g_{d}(n)=O_{d}(n^{2/d})$, and this bound is conjectured to be optimal. The best known lower bound to date is due to Solymosi and Vu [38] which is $g_{d}(n)=\Omega_{d}(n^{2/d-2/(d(d+2))})$.
A well-studied variant of this problem asks for the minimum number $g^{\prime}_{d}(n)$ of distinct $d$-dimensional volumes of $d$-simplices spanned by $n$ points in $\mathbb{R}^{d}$, assuming that not all $n$ points lie on a common hyperplane. One easily sees that $g^{\prime}_{d}(n)\leq\lfloor(n-1)/d\rfloor$ by taking $d$ sets of about $n/d$ equally spaced points on $d$ parallel lines passing through the vertices of a fixed (d1)(d-1)-simplex. Erdős, Purdy, and Straus [16] conjectured that this construction is tight for all sufficiently large $n$. Pinchasi [33] proved this conjecture for the case $d=2$, and Dumitrescu and Cs.Tóth [11] later showed that $g^{\prime}_{d}(n)=\Theta_{d}(n)$ for every $d$.
There is another related problem asking (A): what is the maximum number $h_{d}(n)$ such that any set of $n$ points in $\mathbb{R}^{d}$ contains a subset $S$ of size $h_{d}(n)$ whose induced $d$-simplices all have distinct $d$-volumes. The upper bound is again derived from the $n^{1/d}\times\cdots\times n^{1/d}$ grid, which gives $h_{2}(n)=O_{d}(n^{2/3})$ and $h_{d}(n)=O_{d}(n^{(d-1)/d})$ for $d\geq 3$. Conlon et al. [7] showed $h_{d}(n)=\Omega_{d}(n^{1/(2d+2)})$ for all $d\geq 2$.
@@ 1.2 Erdős’s unit distance problem
The unit distances problem asks: what is the maximum number $u_{d}(n)$ of unit distances that a set of $n$ points in $\mathbb{R}^{d}$ can form. For the case $d=2$, Erdős [12] obtained the bounds $n^{1+c/\log\log n}<u_{2}(n)=O(n^{3/2})$ and conjectured that his lower bound is tight. The problem remains wide open; the best known upper bound $u_{2}(n)=O(n^{4/3})$ is due to Spencer, Szemerédi, and Trotter [39]. Despite numerous efforts [5, 40, 32], no further improvements have been found. In $3$-dimensions, the best known bounds are $u_{3}(n)=O(n^{3/2})$ [26, 44] and $u_{3}(n)=\Omega(n^{4/3}\log\log n)$ [13]. For $d\geq 4$, we have $u_{d}(n)=\Theta_{d}(n^{2})$ (see [30, 14]).
This problem also has a well-studied direction of generalization. In 1967, Oppenheim asked (B): what is the maximum number of unit area triangles determined by $n$ points in the plane? A simple lattice construction $\sqrt{n}\times\sqrt{n}$ again gives a lower bound $\Omega(n^{2}\log\log n)$ [15]. Currently, the best known upper bound is $O(n^{2+2/9})$, due to Raz and Sharir [34]. Dumitrescu and Cs.Tóth [11] extended this problem to three dimensions and obtained the following result: the number of unit volume tetrahedra ($3$-simplices) determined by $n$ points in $\mathbb{R}^{3}$ is $O(n^{7/2})$, and there is an $n$-point set achieving $\Omega(n^{3}\log\log n)$.
Related problems (C): are to determine the number of triangles of minimum/maximum area determined by a set of $n$ points studied by Braß, Rote, and Swanepoel [2]. They proved that the maximum number of triangles of maximum area determined by $n$ points in the plane is exactly $n$, and the maximum number of triangles of minimum area determined by $n$ points in the plane is $\Theta(n^{2})$.
Dumitrescu and Cs.Tóth [11] further generalized this problem and obtained the following results: The number of tetrahedra of minimum (non-zero) volume spanned by $n$ points in $\mathbb{R}^{3}$ is at most $2n^{3}/3-O(n^{2})$, and there is a configulation of $n$-points achieving $3n^{3}/16-O(n^{2})$. Furthermore, they proved that the maximum number of $d$-simplices of minimum volume spanned by $n$ points in $\mathbb{R}^{d}$ is $O_{d}(n^{d})$ for all $d$.
@@ 1.3 Dual setting
    What is the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct volumes?
    For the case $d=2$, Damásdi et al. [9] obtained a lower bound of $n^{1/5}$ under the stronger assumption that there are no six lines that are tangent to a common conic (note that five lines always have a common tangent conic).
@@
    The problem of generalizing (B) to higher dimensions is to determine the maximum number of unit volume $d$-simplices formed by n points in $\mathbb{R}^{d}$. For this problem, we consider the following dual version.
@@
    What is the maximum number $f_{d}(n)$ of unit-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$?
    For the case $d=2$, Damásdi et al. [9] proved that $f_{2}(n)=O(n^{9/4+\varepsilon})$ for every fixed $\varepsilon>0$, and $f_{2}(n)=\Omega(n^{2})$. Their proof uses the incidence theorem between points and algebraic curves, proved by Sharir and Zahl [36].
@@
    We propose the following variants for the higher-dimensional generalization of (B) (that is, to determinie the number of $d$-simplices of minimum/maximum volume determined by a set of $n$ points in $\mathbb{R}^{d}$).
@@
    What is the maximum number $m_{d}(n)$ of minimum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    For the case $d=2$, Damásdi et al. [9] proved that $m_{2}(n)=\Theta(n^{2})$.
@@
    What is the maximum number $M_{d}(n)$ of maximum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    As we mentioned above, with respect to the occurrence of the maximum area determined by $n$ points in the plane, it is known that the maximum number is exactly $n$. However, Damásdi et al. [9] showed that this phenomenon does not occur in the dual setting: $M_{2}(n)>7n/5-O(1)$. They also showed an upper bound $M_{2}(n)\leq n(n-1)/3$.
@@ 1.4 Our results
For the maximum number $f_{d}(n)$ of unit-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
@@
For the maximum number $m_{d}(n)$ of minimum-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
@@
For the maximum number $M_{3}(n)$ of maximum-volume tetrahedra formed by $n$ planes in $\mathbb{R}^{3}$, we have
$M_{3}(n)>\frac{7}{6}n-O(1).$
@@
$D_{2}(n)<<n(\log n)^{-c}$ for some absolute constant $c>0$, and for $d\geq 3$, $D_{d}(n)<<ne^{-(\log\log n)^{c_{d}}}$ for some absolute constant $c_{d}\in(0,1)$.
Theorem 1.5 gives the a partial answer for Problem 5.4 and Problem 5.5 in [9].
@@ 2 Unit volume $d$-simplices
In this section, we consider the number of $d$-simplices with unit volume determined by an arrangement of $n$ hyperplanes, namely to investigate the magnitude of $f_{d}(n)$. We have a trivial upper bound $O_{d}(n^{d+1})$ by selecting generic $d+1$ hyperplanes from the arrangement. The following theorem provides a slightly better upper bound for $f_{d}(n)$. See 1.1 First, we see the following proposition.
@@
Let $H_{1},\ldots,H_{d}$ be hyperplanes in $\mathbb{R}^{d}$ whose intersection is a single point. For a fixed coordinate system $(x_{1},\ldots,x_{d})$, there exist exactly two hypersurfaces of degree $d$ obtained by an (invertible) affine transformation of the hypersurface $\prod_{i=1}^{d}x_{i}=c$ for some $c\neq 0$ such that each hyperplane tangent to them and $H_{1},\ldots,H_{d}$ form a $d$-simplex with unit volume.
@@
Note that an (invertible) affine transformations preserve hyperplanes, hypersurfaces of degree $d$ and ratios of $d$-volumes. Therefore, by applying an appropriate affine transformation, we may set $H_{i}=\{x_{i}=0\}$ for $i\in\{1,\ldots,d\}$. Now we fix $c>0$ and consider the hypersurface $\prod_{i=1}^{d}x_{i}=c$. Then the equation of the hyperplane $h_{p}$ tangent to this surface at a point $p=(p_{1},\ldots,p_{d})$ is given by $\sum_{i=1}^{d}x_{i}/p_{i}=d$. The hyperplane $h_{p}$ and $H_{1},\ldots,H_{d}$ form a $d$-simplex $\Delta_{p}$ with vertices $(0,\ldots,0,dp_{i},0,\ldots,0)$ $(1\leq i\leq d)$, where $dp_{i}$ is in the $i$-th coordinate. Hence, the volume of $\Delta_{p}$ is
$\frac{1}{d!}\cdot\prod_{i=1}^{d}dp_{i}=\frac{d^{d}}{d!}|c|,$
which is a constant depending only on the dimension $d$. Similarly for $c<0$, any tangent hyperplane of $x_{1}x_{2}\cdots x_{d}=c$ and $H_{1},\ldots,H_{d}$ form a $d$-simplex with the same volume $d^{d}|c|/d!$. ∎
@@
The proof is based on [10]. Let $\mathcal{H}$ be an arrangement of $n$ hyperplanes. Each of the two hypersurfaces given in the above proposition has $2^{d-1}$ connected branches ($2^{d}$ connected branches in total). Let $\sum\nolimits_{d}$ be a set of all branches obtained by selecting $d$ hyperplanes intersecting at a single point from $\mathcal{H}$. Then, we have $|\sum\nolimits_{d}|\leq 2^{d}\binom{n}{d}$. Note that for each branch $\sigma\in\sum\nolimits_{d}$, the dual of the set of all hyperplanes tangent to $\sigma$ is a subset of some hypersurface of degree $d$, and the number of its connected component is: two when there exists a hyperplane tangent to $\sigma$ parallel to the last axis, and one otherwise. By dual transformation in $\mathbb{R}^{d}$, we mean the bijection between points $p=(p_{1},p_{2},\ldots,p_{d-1},p_{d})\in\mathbb{R}^{d}$ and hyperplanes $p^{*}:x_{d}=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{d-1}x_{d-1}+p_{d}$ in $\mathbb{R}^{d}$. Now we consider two bipartite graphs $G_{1},G_{2}$ with vertex set $\mathcal{H}\times\sum\nolimits_{d}$ as follows. If $H\in\mathcal{H}$ is tangent to $\sigma\in\sum\nolimits_{d}$ from above (in the sense of the last axis), we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{1}$. Similarly, If $H\in\mathcal{H}$ is tangent to $S\in\sum\nolimits_{d}$ from below, we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{2}$. The edges of $G_{1}$ and $G_{2}$ correspond to all hyperplane-hypersurface tangencies. Thus, to obtain an upper bound for $f_{d}(n)$, it suffices to bound the number of edges $E(G_{1})$ in $G_{1}$ and the number of edges $E(G_{2})$ in $G_{2}$. Observe that neither $G_{1}$ nor $G_{2}$ contains $K_{d^{d}+1,d}$. Indeed, if $G_{1}$ contains $K_{d^{d}+1,d}$, then there would be $d^{d}+1$ hyperplanes in $\mathcal{H}$ tangent to $d$ branches from above. Considering this situation in a dual space, meaning that $d$ branches intersect at $d^{d}+1$ points, which contradicts the higher-dimensional Bézout’s theorem (refer to [42]). The same argument applies to $G_{2}$. Therefore, by the Kövári–Sós–Turán theorem,
$|E(G_{1})|,|E(G_{2})|=O\bigl(n\cdot n^{d(1-\frac{1}{d^{d}+1})}+n^{d}\bigr)=O\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
Clearly, a lower bound for $m_{d}(n)$ gives a lower bound for $f_{d}(n)$. See the result in the next section.

### v2

@@ Abstract
The distinct distances problem and the unit distance problem, posed by Erdős [erdos1946] in 1946, are classical problems in combinatorial geometry. They ask for the minimum number of distinct distances determined by $n$ points in $\mathbb{R}^{d}$ and the maximum number of pairs at unit-distance that such a set can determine, respectively. These problems have been studied extensively for several decades, leading to numerous deep results. We study the following dual setting:
  • Determine the maximum number of unit $d$-dimensional $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number of minimum (or maximum) $d$-dimensional $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$.
  • Determine the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ for which all induced $d$-dimensional simplices have distinct $d$-dimensional volumes.
@@ 1 Introduction
For $d\geq 2$, a finite point set $X$ in $\mathbb{R}^{d}$ with $|X|\geq d+1$ is said to be in general position if no $d+1$ points of $X$ lie on the same hyperplane, and we say that an arrangement of hyperplanes $\mathcal{A}$ in $\mathbb{R}^{d}$ with $|\mathcal{A}|\geq d+1$ is in general position if every $d$ hyperplanes intersect at exactly one point and no $d+1$ hyperplanes have a common point.
Throughout the paper, we use $O_{d}$ to denote the standard big-$O$ notation in which the constant may depend on $d$. We use analogous conventions for $\Omega_{d}$ and $\Theta_{d}$. For $1\leq k\leq d$, we will refer to $k$-dimensional volume and $k$-dimensional simplex as $k$-volume and $k$-simplex. Furthermore, by “a $k$-simplex in a hyperplane arrangement in $\mathbb{R}^{d}$”, we mean a simplicial $k$-cell determined by $k+1$ hyperplanes of the arrangement.
1.1 Erdős’s distinct distances problem
In 1946, Erdős [erdos1946] posed the question: what is the minimum number $g_{d}(n)$ of distinct distances determined by $n$ points in $\mathbb{R}^{d}$. This question is known as Erdős’s distinct distances problem, and a variety of interesting results have been obtained. For the planar case $d=2$, Erdős showed in the same paper an upper bound $g_{2}(n)=O(n/\sqrt{\log n})$ by considering the $\sqrt{n}\times\sqrt{n}$ square lattice. While there had been earlier progress on lower bounds ([chung1984number, chung1992number, katz2004new, moser1952different, solymosi2001distinct, ekely1993crossing, tardos2003distinct]), a major breakthrough by Guth and Katz [guth2015erdHos] proved $g_{2}(n)=\Omega(n/\log n)$.
In higher dimensions, a $d$-dimensional grid $n^{1/d}\times\cdots\times n^{1/d}$ gives $g_{d}(n)=O_{d}(n^{2/d})$, and this bound is conjectured to be optimal. The best known lower bound to date is due to Solymosi and Vu [solymosi2008near] which is $g_{d}(n)=\Omega_{d}(n^{2/d-2/(d(d+2))})$.
A well-studied variant of this problem asks for the minimum number $g^{\prime}_{d}(n)$ of distinct $d$-volumes of $d$-simplices spanned by $n$ points in $\mathbb{R}^{d}$, assuming that not all $n$ points lie on a common hyperplane. One easily sees that $g^{\prime}_{d}(n)\leq\lfloor(n-1)/d\rfloor$ by taking $d$ sets of about $n/d$ equally spaced points on $d$ parallel lines passing through the vertices of a fixed (d1)(d-1)-simplex. Erdős, Purdy, and Straus [erdos1982problem] conjectured that this construction is tight for all sufficiently large $n$. Pinchasi [pinchasi2008minimum] proved this conjecture for the case $d=2$, and Dumitrescu and Cs.Tóth [dumitrescu2008number] later showed that $g^{\prime}_{d}(n)=\Theta_{d}(n)$ for every $d$.
There is another related problem asking (A): what is the maximum number $h_{d}(n)$ such that any set of $n$ points in $\mathbb{R}^{d}$ contains a subset $S$ of size $h_{d}(n)$ whose induced $d$-simplices all have distinct $d$-volumes. The upper bound is again derived from the $n^{1/d}\times\cdots\times n^{1/d}$ grid, which gives $h_{2}(n)=O_{d}(n^{2/3})$ and $h_{d}(n)=O_{d}(n^{(d-1)/d})$ for $d\geq 3$. Conlon et al. [conlon2015distinct] showed $h_{d}(n)=\Omega_{d}(n^{1/(2d+2)})$ for all $d\geq 2$.
@@ 1.2 Erdős’s unit distance problem
The unit distance problem asks: what is the maximum number $u_{d}(n)$ of unit distances that a set of $n$ points in $\mathbb{R}^{d}$ can form. For the case $d=2$, Erdős [erdos1946] obtained the bounds $n^{1+c/\log\log n}<u_{2}(n)=O(n^{3/2})$ for some $c>0$ and conjectured that his lower bound is tight. The problem remains wide open; the best known upper bound $u_{2}(n)=O(n^{4/3})$ is due to Spencer, Szemerédi, and Trotter [spencer1984unit]. Despite numerous efforts ([clarkson1990combinatorial, pach2006forbidden, ekely1993crossing]), no further improvements have been found. In $3$-dimensions, the best known bounds are $u_{3}(n)=O(n^{3/2})$ [kaplan2012unit, zahl2013improved] and $u_{3}(n)=\Omega(n^{4/3}\log\log n)$ [erdos1960set]. For $d\geq 4$, we have $u_{d}(n)=\Theta_{d}(n^{2})$ (see [erdHos1990variations, lenz1955zerlegung]).
This problem also has a well-studied direction of generalization. In 1967, Oppenheim asked (B): what is the maximum number of unit area triangles determined by $n$ points in the plane? See [erdHos1996extremal]. A simple lattice $\sqrt{\log n}\times(n/\sqrt{\log n})$ again gives a lower bound $\Omega(n^{2}\log\log n)$ [erdos1971some]. Currently, the best known upper bound is $O(n^{2+2/9})$, due to Raz and Sharir [raz2017number]. Dumitrescu and Cs.Tóth [dumitrescu2008number] extended this problem to three dimensions and obtained the following result: the number of unit volume tetrahedra ($3$-simplices) determined by $n$ points in $\mathbb{R}^{3}$ is $O(n^{7/2})$, and there is an $n$-point set achieving $\Omega(n^{3}\log\log n)$.
Related problems (C): are to determine the number of triangles of minimum/maximum area determined by a set of $n$ points studied by Braß, Rote, and Swanepoel [brass2001triangles]. They proved that the maximum number of triangles of maximum area determined by $n$ points in the plane is exactly $n$, and the maximum number of triangles of minimum area determined by $n$ points in the plane is $\Theta(n^{2})$.
Dumitrescu and Cs.Tóth [dumitrescu2008number] further generalized this problem and obtained the following results: The number of tetrahedra of minimum volume spanned by $n$ points in $\mathbb{R}^{3}$ is at most $2n^{3}/3-O(n^{2})$, and there is a configuration of $n$-points achieving $3n^{3}/16-O(n^{2})$. Furthermore, they proved that the maximum number of $d$-simplices of minimum volume spanned by $n$ points in $\mathbb{R}^{d}$ is $O_{d}(n^{d})$ for all $d$.
@@ 1.3 Dual setting
    What is the maximum number $D_{d}(n)$ such that any arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$ in general position contains a subset of size $D_{d}(n)$ whose induced $d$-simplices have all distinct $d$-volumes?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] obtained a lower bound of $n^{1/5}$ under the stronger assumption that there are no six lines that are tangent to a common conic (note that five lines always have a common tangent conic).
@@
    The problem of generalizing (B) to higher dimensions is to determine the maximum number of unit $d$-volume $d$-simplices formed by $n$ points in $\mathbb{R}^{d}$. For this problem, we consider the following dual version.
@@
    What is the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] proved that $f_{2}(n)=O(n^{9/4+\varepsilon})$ for every fixed $\varepsilon>0$, and $f_{2}(n)=\Omega(n^{2})$. Their proof uses the incidence theorem between points and algebraic curves, proved by Sharir and Zahl [sharir2017cutting].
@@
    We propose the following variants for the higher-dimensional generalization of (C) (that is, to determinie the maximum number of $d$-simplices of minimum/maximum $d$-volume determined by a set of $n$ points in $\mathbb{R}^{d}$).
@@
    What is the maximum number $m_{d}(n)$ of minimum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    For the case $d=2$, Damásdi et al. [damasdi2020triangle] proved that $m_{2}(n)=\Theta(n^{2})$.
@@
    What is the maximum number $M_{d}(n)$ of maximum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$ ?
    As we mentioned above, with respect to the occurrence of the maximum area determined by $n$ points in the plane, it is known that the maximum number is exactly $n$. However, Damásdi et al. [damasdi2020triangle] showed that this phenomenon does not occur in the dual setting: $M_{2}(n)>7n/5-O(1)$. We note that they also showed an upper bound $M_{2}(n)\leq n(n-1)/3$.
@@ 1.4 Our results
For the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
@@
For the maximum number $f_{d}(n)$ of unit $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
@@
For the maximum number $m_{d}(n)$ of minimum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
@@
For the maximum number $M_{d}(n)$ of maximum $d$-volume $d$-simplices formed by $n$ hyperplanes in $\mathbb{R}^{d}$, we have
$M_{d}(n)>\frac{7}{5}(n-d)-O(1).$
@@
For the largest subset of hyperplanes forming $d$-simplices of distinct $d$-volumes, we have
$D_{2}(n)\ll n(\log n)^{-c}$
for some absolute constant $c>0$, For every $d\geq 3$, there is $c_{d}\in(0,1)$ such that
$D_{d}(n)\ll ne^{-(\log\log n)^{c_{d}}}.$
Theorem 1.5 gives partial answers for Problem 5.4 and Problem 5.5 in [damasdi2020triangle].
@@ 2 Unit volume $d$-simplices
In this section, we consider the number of $d$-simplices with unit volume determined by an arrangement of $n$ hyperplanes, namely to investigate the magnitude of $f_{d}(n)$. We have a trivial upper bound $O_{d}(n^{d+1})$ by selecting generic $d+1$ hyperplanes from the arrangement. The following provides a slightly better upper bound for $f_{d}(n)$. See 1.1
@@
Let $H_{1},\dots,H_{d}$ be hyperplanes in $\mathbb{R}^{d}$ meeting at a unique point and fix $V_{0}>0$. Then, every hyperplane that form a $d$-simplex with $H_{1},\ldots,H_{d}$ of $d$-volume $V_{0}$ are all tangent to one of two fixed algebraic hypersurfaces of degree $d$ which have $H_{1},\ldots,H_{d}$ as asymptotes and each has exactly $2^{d-1}$ connected components. Moreover, the $2^{d}$ connected components of these two hypersurfaces are contained one each in the $2^{d}$ connected components determined by $H_{1},\ldots,H_{d}$.
@@
Note that an (invertible) affine transformation preserves hyperplanes, hypersurfaces of degree $d$ and ratios of volumes. Fix a coordinate $(x_{1},\ldots,x_{d})$ and we may assume $H_{i}=\{x_{i}=0\}\quad(i=1,\ldots,d).$ Fix $c\neq 0$ and consider the hypersurface
$S_{c}:=\Bigl\{x\in\mathbb{R}^{d}:\ \prod_{i=1}^{d}x_{i}=c\Bigr\}.$
Take an arbitrary point $p=(p_{1},\dots,p_{d})\in S_{c}$. Then $\prod_{i=1}^{d}p_{i}=c$, and $p_{i}\neq 0$ for all $i$. The tangent hyperplane of $S_{c}$ at $p$ is given by $\sum_{i=1}^{d}\frac{x_{i}}{p_{i}}=d.$ This tangent hyperplane together with $H_{1},\dots,H_{d}$ forms a $d$-simplex $\Delta_{p}$ with vertices
$O=(0,\dots,0),\quad(dp_{1},0,\dots,0),\quad(0,dp_{2},0,\dots,0),\ \dots,\ (0,\dots,0,dp_{d}).$
Hence the volume of $\Delta_{p}$ equals
$\frac{1}{d!}\prod_{i=1}^{d}|dp_{i}|=\frac{d^{d}}{d!}\,\Bigl|\prod_{i=1}^{d}p_{i}\Bigr|=\frac{d^{d}}{d!}\,|c|.$
Since the volume of $\Delta_{p}$ depends only on $d$ and $|c|$, for any point of $S_{-c}$, the tangent hyperplane at that point together with $H_{1},\ldots,H_{d}$ also cuts out a $d$-simplex of volume $\frac{d^{d}}{d!}\,|c|$. The hypersurfaces $S_{c}$ and $S_{-c}$ have $2^{d-1}$ connected components each, and every connected components has $H_{1},\ldots,H_{d}$ as asymptotes. Moreover, each of the $2^{d}$ regions determined by $H_{1},\ldots,H_{d}$ contains exactly one connected component of $S_{c}\cup S_{-c}$. Consequently, any hyperplane that, together with $H_{1},\ldots,H_{d}$, forms a $d$-simplex of volume $\frac{d^{d}}{d!}\,|c|$ is tangent to exactly one connected component of $S_{c}\cup S_{-c}$. One easily sees that the $2^{d}$ connected components of $S_{c}\cup S_{-c}$ are contained one each in the $2^{d}$ connected components determined by $H_{1},\ldots,H_{d}$. ∎
@@
The proof is based on [dumitrescu2009extremal]. Let $\mathcal{H}$ be an arrangement of $n$ hyperplanes in $\mathbb{R}^{d}$. For every $d$-tuple $(H_{i_{1}},\ldots,H_{i_{d}})$ of hyperplanes in $\mathcal{H}$ with $\bigcap_{j=1}^{d}H_{i_{j}}$ consisting of a single point, apply the above Proposition with $V_{0}=1$ to obtain the two algebraic hypersurfaces of degree $d$. Let $\sum\nolimits_{d}$ be the union of all such connected components by choosing $d$ hyperplanes intersecting at a single point. Since each hypersurface has $2^{d-1}$ connected components, we have $|\sum\nolimits_{d}|\leq 2\cdot 2^{d-1}\binom{n}{d}$.
We consider the following dual transformation in $\mathbb{R}^{d}$ between hyperplanes $H:x_{d}=p_{1}x_{1}+p_{2}x_{2}+\cdots+p_{d-1}x_{d-1}+p_{d}$ and points $H^{*}=(p_{1},p_{2},\ldots,p_{d-1},p_{d})$. For a connected component $\sigma\in\sum\nolimits_{d}$, define $X_{\sigma}:=\{H^{*}:H\text{ is tangent to }\sigma\}.$ Then, $X_{\sigma}$ is contained in an algebraic hypersurface of degree at most $d$. Moreover, $X_{\sigma}$ has at most two connected components: it has two connected components if and only if $\sigma$ admits a tangent hyperplane parallel to the $x_{d}$-axis.
Now we consider two bipartite graphs $G_{1},G_{2}$ with vertex set $\mathcal{H}\sqcup\sum\nolimits_{d}$ as follows. If $H\in\mathcal{H}$ is tangent to $\sigma\in\sum\nolimits_{d}$ from above (in the sense of the last axis), we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{1}$. Similarly, If $H\in\mathcal{H}$ is tangent to $S\in\sum\nolimits_{d}$ from below, we define an edge $(H,\sigma)\in\mathcal{H}\times\sum\nolimits_{d}$ in $G_{2}$. The edges of $G_{1}$ and $G_{2}$ correspond to all hyperplane-hypersurface tangencies. Hence, to obtain an upper bound for $f_{d}(n)$, it suffices to bound the number of edges $E(G_{1})$ in $G_{1}$ and the number of edges $E(G_{2})$ in $G_{2}$. Observe that neither $G_{1}$ nor $G_{2}$ contains $K_{d^{d}+1,d}$ as a subgraph. Indeed, if $G_{1}$ contains $K_{d^{d}+1,d}$, then there would be $d^{d}+1$ hyperplanes in $\mathcal{H}$ tangent to $d$ connected components from above. Considering this situation in a dual space, meaning that $d$ connected components intersect at $d^{d}+1$ points, which contradicts the higher-dimensional Bézout’s theorem (refer to [tao2011bezout]). The same argument applies to $G_{2}$. Therefore, by the Kövári–Sós–Turán theorem,
$|E(G_{1})|,|E(G_{2})|=O_{d}\bigl(n\cdot n^{d(1-\frac{1}{d^{d}+1})}+n^{d}\bigr)=O_{d}\bigl(n^{\,d+1-\tfrac{d}{d^{d}+1}}\bigr).$
It is clear that a lower bound for $m_{d}(n)$ gives a lower bound for $f_{d}(n)$. See the lower bound construction in the next section.